it takes a garbage scow 8 hours to travel 24 miles up stream and 8 hours to travel 88 miles down stream. what is average speed of the scow in still water , and what is the speed of the current.
Answer by jorel1380(3719)

24+88/16=112/16=7
24/8=3 mph
88/8=11 mph
7-c=3
7+c=11
c=4
The speed of the boat in still water is 7 mph. The current is 3 mph..